# Capture of field stars by dark substructures

### v1

@@ 5 Summary
Substructures orbiting a larger galaxy can capture field stars that pass nearby with low relative velocities. In this work, stars become captured by the substructure potential $\Phi_{\bullet}$ when their gravitational energy $E=v^{2}/2+\Phi_{\bullet}(r)$ flips from positive to negative.
Our numerical experiments show that the orbits of stars captured through this mechanism can be separated into two families: (i) Temporary captures. The majority of field stars trapped in the potential $\Phi_{\bullet}$ move on chaotic orbits that orbit around the substructure for a finite amount of time before being released back to the galactic potential. This leads to a net number of temporary captures that converges to a steady state as the rate of capture becomes roughly equal to the escape rate. Stars captured on temporary orbits generate stellar overdensities ($\delta_{\star}>1)$ at the location of substructures that are sufficiently compact, $\kappa=1-c_{\bullet}/r_{\epsilon}>0$, where $r_{\epsilon}$ and $c_{\bullet}$ are the thermal critical radius (6) and scale radius of $\Phi_{\bullet}$, respectively. (ii) Permanent captures. In addition, we find that immersing a substructure potential in a galaxy that is in dynamical equilibrium at $t=0$ invariably results in a population of field particles that remain bound to $\Phi_{\bullet}$ for arbitrarily-long times. In static models where $\Phi_{\bullet}$ does not vary with time and moves on a circular orbit around the host galaxy, we find that particles on permanent orbits were already bound to the substructure potential at $t=0$, and remain within its tidal radius indefinitely. These particles are therefore reminiscent of the ‘potential escapers’ discussed by Fukushige & Heggie (2000). Our experiments also show that substructures with a time-growing mass, $M_{\bullet}(t)$, can also capture stars on permanent orbits at $t>0$, with a rate that peaks at $t\sim\tau$, where $\tau=|\dot{M}_{\bullet}/M_{\bullet}|^{-1}_{t=0}$ is the time-scale that determines how fast the substructure mass varies. Importantly, we find that stars on permanent orbits have a density that lies below the host galaxy background ($\delta_{\star}<1$), and therefore do not generate visible overdensities.
@@
We have given numerical and theoretical arguments that indicate that ‘dark’ subhaloes that were not sufficiently massive to trigger in-situ star formation may become ‘visible’ by capturing baryonic particles from the host galaxy field. The important implication of this result is that dark subhaloes cannot be completely invisible. If they contain gravitationally-bound baryonic matter, they must emit and absorb radiation, which opens up a new avenue to test CDM predictions on halo mass scales that have not been probed to date. This calls for observational campaigns aimed at detecting & characterizing ‘dark’ subhaloes in dSphs with masses below the star formation threshold. In §4.3 we discuss how the detection and characterization of these objects may provide unique constraints on the particle mass and cross section for a large range of DM candidates.
@@ Acknowledgements
We thank Sergey Koposov, Alan McConnachie and Frank van den Bosch for their valuable comments and suggestions. M.G.W. acknowledges support from National Science Foundation (NSF) grants AST-1813881, AST-1909584 and AST-2206046. M.G. acknowledges financial support from the grants PID2021-125485NB-C22, EUR2020-112157, CEX2019-000918-M funded by MCIN/AEI/10.13039/501100011033 (State Agency for Research of the Spanish Ministry of Science and Innovation) and SGR-2021-01069 (AGAUR). RE acknowledges support from the European Research Council (ERC) under the European Unions Horizon 2020 research and innovation programme (grant agreement number 834148), and from the National Science Foundation (NSF) grant AST-2206046. TB’s research was supported by an appointment to the NASA Postdoctoral Program at the NASA Ames Research Center, administered by Oak Ridge Associated Universities under contract with NASA.

### v2

@@ 5 Summary
Substructures orbiting a larger galaxy can capture field stars that pass nearby with low relative velocities. In this work, stars become captured by the substructure potential $\Phi_{\bullet}$ when their gravitational energy $E=v^{2}/2+\Phi_{\bullet}(r)$ flips from positive to negative. Our numerical experiments show that the orbits of stars captured through this mechanism can be separated into two families: (i) Temporary captures. The majority of field stars trapped in the potential $\Phi_{\bullet}$ move on chaotic orbits that orbit around the substructure for a finite amount of time before being released back to the galactic potential. This leads to a net number of temporary captures that converges to a steady state as the rate of capture becomes roughly equal to the escape rate. Stars captured on temporary orbits generate stellar overdensities ($\delta_{\star}>1)$ at the location of substructures that are sufficiently compact, $\kappa=1-c_{\bullet}/r_{\epsilon}>0$, where $r_{\epsilon}$ and $c_{\bullet}$ are the thermal critical radius (9) and scale radius of $\Phi_{\bullet}$, respectively. (ii) Permanent captures. In addition, we find that immersing a substructure potential in a galaxy that is in dynamical equilibrium at $t=0$ invariably results in a population of field particles that remain bound to $\Phi_{\bullet}$ for arbitrarily-long times. In static models where $\Phi_{\bullet}$ does not vary with time and moves on a circular orbit around the host galaxy, we find that particles on permanent orbits were already bound to the substructure potential at $t=0$, and remain within its tidal radius indefinitely. These particles are therefore reminiscent of the ‘potential escapers’ discussed by Fukushige & Heggie (2000). Our experiments also show that substructures with a time-growing mass, $M_{\bullet}(t)$, can also capture stars on permanent orbits at $t>0$, with a rate that peaks at $t\sim\tau$, where $\tau=|\dot{M}_{\bullet}/M_{\bullet}|^{-1}_{t=0}$ is the time-scale that determines how fast the substructure mass varies. Importantly, we find that stars on permanent orbits have a density that lies below the host galaxy background ($\delta_{\star}<1$), and therefore do not generate visible overdensities.
@@
We have given numerical and theoretical arguments that indicate that ‘dark’ subhaloes that were not sufficiently massive to trigger in-situ star formation may become ‘visible’ by capturing baryonic particles from the host galaxy field. The important implication of this result is that dark subhaloes may not be completely invisible. If they contain gravitationally-bound baryonic matter, they must emit and absorb radiation, which opens up a new avenue to test CDM predictions on halo mass scales that have not been probed to date. Follow-up work is needed to inspect the detectability of ‘dark’ subhaloes in dSphs with masses below the star formation threshold, as well as observational campaigns aimed at detecting & characterizing those objects. In §4.3 we discuss how the detection and characterization of dark subhaloes may provide unique & unprecedented constraints on the particle mass and cross section for a large range of DM candidates.
@@ Acknowledgements
We are grateful to the referee for the constructive criticism and the suggestions that helped to improve the paper. We thank Sergey Koposov, Alan McConnachie and Frank van den Bosch for their valuable comments and suggestions. M.G.W. acknowledges support from National Science Foundation (NSF) grants AST-1813881, AST-1909584 and AST-2206046. M.G. acknowledges financial support from the grants PID2021-125485NB-C22, EUR2020-112157, CEX2019-000918-M funded by MCIN/AEI/10.13039/501100011033 (State Agency for Research of the Spanish Ministry of Science and Innovation) and SGR-2021-01069 (AGAUR). RE acknowledges support from the European Research Council (ERC) under the European Unions Horizon 2020 research and innovation programme (grant agreement number 834148), and from the National Science Foundation (NSF) grant AST-2206046. TB’s research was supported by an appointment to the NASA Postdoctoral Program at the NASA Ames Research Center, administered by Oak Ridge Associated Universities under contract with NASA.
In memory of Teo Peñarrubia Reihl (16-3-2018, 6-1-2024). V.A.E.J.H.